OpenStudy (anonymous):

Keisha's parents want to save $20,000 in her college savings account over the next 15 years. They have $8,000 to use as an initial deposit. What simple annual interest rate do they need to meet their goal? Round your answer to the nearest tenth.

OpenStudy (anonymous):

\[SI =\frac{ P \times R \times T }{ 100 }\]

OpenStudy (anonymous):

SI=Simple Interest P=Principal (Initial deposit) R=Rate T=Time

OpenStudy (anonymous):

can you continue from there???

OpenStudy (anonymous):

Um, let me try.

OpenStudy (anonymous):

What is the rate?

OpenStudy (anonymous):

\[20,000 = 1200R\]

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

16.6%

OpenStudy (anonymous):

That isn't right.

OpenStudy (anonymous):

Err, I don't know. Can you help me, are you still there?